# Closed-Loop Neural Operator-Based Observer of Traffic Density

### v1

@@ II-C Problem Statement
Following this discussion, we model the SUMO simulations as an $n^{\text{th}}$-order discrete-time dynamical system
$\mathcal{S}_{\text{SUMO}}[\bar{\zeta_{0}}]:\begin{cases}\rho(\cdot,t+\Delta t)=\mathcal{G}[\zeta_{t}],\quad t>0,\\ \zeta_{t}=[\rho(\cdot,t),\rho(\cdot,t-\Delta t),\dots],\\ \zeta_{0}=\bar{\zeta_{0}},\end{cases}$ (5)
where $\rho(\cdot,t)\in\mathcal{H}^{1}_{\Omega}$ is the true density, $\zeta_{t}\in\mathcal{H}^{n}_{\Omega}$ is the state at time $t$, $\bar{\zeta_{0}}$ is the initial state, and $\mathcal{G}$ represents the solution operator of the unknown dynamics.
The objective of this paper is to develop an algorithm in the form of a dynamical system that observes (5) in closed-loop by generating real-time predictions $\hat{\rho}$ of the system state $\rho$ every $\Delta t$ time unit such that $\|\rho-\hat{\rho}\|$ is minimized. To achieve this goal, we assume access to online measurements $y(x,t)=\rho(x,t)+\epsilon$ with i.i.d. noise $\epsilon$ at stationary, sparse sensor locations $x\in\mathcal{M}\subset\Omega$, collectively denoted $\mathbf{y}_{\mathcal{M}}(t)$. We also assume access to offline data from physically identical systems, $\{\mathcal{S}_{\text{SUMO}}[\zeta_{0}^{i}]\}_{i\in\mathcal{I}}$.
@@ III Methodology
To build a closed-loop observer of (5), the first step is to obtain a representation of the open-loop prediction, $\mathcal{G}$. Thus, our approach starts by learning an approximate solution operator $\mathcal{G}_{\theta}\approx\mathcal{G}$ from offline SUMO data. We then compare two methods of iteratively applying $\mathcal{G}$ to observe (5) in an open-loop manner, $\hat{\mathcal{S}}^{ol}$ and $\hat{\mathcal{S}}^{ol\text{-}r}$. To close the loop, we finally design an algorithm, $\hat{\mathcal{S}}^{cl}$, that integrates sparse online measurements with recursive predictions using $\mathcal{G}$. The research problem is then addressed by using $\hat{\mathcal{S}}^{cl}$ together with $\mathcal{G}_{\theta}$, resulting in a data-driven, closed-loop observer.
@@ III-A Solution Operator: Model Identification
Given the offline SUMO scenarios $\{\mathcal{S}_{\text{SUMO}}[\zeta_{0}^{i}]\}_{i\in\mathcal{I}}$, the goal is to learn a representation of the solution operator
$\mathcal{G}_{k}:\begin{array}[t]{rcl}\mathcal{H}^{n}_{\Omega}&\to&\mathcal{H}^{1}_{\Omega},\\ \zeta_{0}&\mapsto&\rho_{k\Delta t},\end{array}$
which maps the input state $\zeta_{0}$ to the solution function at $k$ time steps later, $\rho_{k\Delta t}$. While both $\zeta_{0}$ and $\rho_{k\Delta t}$ are infinite-dimensional functions, SUMO outputs finite-dimensional representations by evaluating them on a uniformly discretized grid $\mathcal{X}\subset\Omega$. This yields $\zeta_{0}\rvert_{\mathcal{X}}\in\mathbb{R}^{n\times\lvert\mathcal{X}\rvert}$ and $\rho_{k\Delta t}\rvert_{\mathcal{X}}\in\mathbb{R}^{\lvert\mathcal{X}\rvert}$. To bridge the gap between finite-dimensional data and function space learning, the approximate solution operator $\mathcal{G}_{\theta}$ is modeled as an FNO.
@@
$\mathcal{G}_{\theta}:\begin{array}[t]{rcl}\mathcal{H}^{n}_{\Omega}&\to&\mathcal{H}^{n_{out}}_{\Omega},\\ \zeta_{0}&\mapsto&[\hat{\rho}_{\Delta t},\hat{\rho}_{2\Delta t},\dots,\hat{\rho}_{n_{out}\Delta t}].\end{array}$ (6)
Including the horizon up to the target estimate $\hat{\rho}_{n_{out}\Delta t}$ was empirically found to improve performance, and it reflects the setup in [20]. Thus, the corresponding learning objective is
[…]
Fig. 1 illustrates a forward pass of a learned operator $\mathcal{G}_{\theta}$ for $n_{out}=100$. We refer to this as open-loop prediction in the observer framework.
Figure 1: Open-loop prediction with solution operator $\mathcal{G}_{\theta}$. It predicts $\hat{\rho}_{\Delta t},\hat{\rho}_{2\Delta t},...,\hat{\rho}_{n_{out}\Delta t}$ given an initial state $\zeta=[\rho_{0},\rho_{-\Delta t},\dots]$. Here, $n_{out}=100$.
@@ III-B Open-Loop Observers
$\hat{\zeta}_{t}=[\hat{\rho}(\cdot,t),\hat{\rho}(\cdot,t-\Delta t),\dots]\in\mathcal{H}^{n}_{\Omega}.$
Thus, triggering an autoregressive rollout only requires an estimate of the initial state $\hat{\zeta}_{0}$. However, this is only partially observed in our problem, since measurements are restricted to the sparse sensor locations $\mathcal{M}\subsetneq\Omega$. To complete the initial state estimate, we interpolate between the sparse measurements $\mathbf{y}_{\mathcal{M}}(t)$ with GP regression,
$\hat{\zeta}_{0}\sim\mathcal{GP}_{[\mathbf{y}_{\mathcal{M}}(0),\mathbf{y}_{\mathcal{M}}(-\Delta t),\dots]}\in\mathcal{H}^{n}_{\Omega}.$
@@
We found that using multi-step ahead prediction $\mathcal{G}_{n_{d}+1}$ with a delayed state estimate $\hat{\zeta}_{t-n_{d}\Delta t}$ offset by $n_{d}>n-1$ improved stability. This creates the open-loop observer $\hat{\mathcal{S}}^{ol}$:
@@
$\hat{\mathcal{S}}^{ol}[\mathcal{G},\mathbf{y}_{\mathcal{M}}]:\left\{\begin{array}[]{l}\hat{\rho}(\cdot,t+\Delta t)=\mathcal{G}_{n_{d}+1}[\hat{\zeta}_{t-n_{d}\Delta t}],\\ \hfill t=0,\Delta t,\dots\\ \hat{\zeta}_{t-n_{d}\Delta t}=[\hat{\rho}(\cdot,t-n_{d}\Delta t),\dots],\\ \hat{\zeta}_{-n_{d}\Delta t}\sim\mathcal{GP}_{[\mathbf{y}_{\mathcal{M}}(-n_{d}\Delta t),\dots]}.\end{array}\right.$ (12)
@@
The gap between the initial state and the first estimate, $\hat{\rho}(\cdot,t),$ $t\in[-n_{d}\Delta t+\Delta t,\dots,0]$, is filled by interpolating the corresponding measurements with GP regression. The complete algorithm is outlined in Algorithm 1.
input : $\mathcal{G},\mathbf{y}_{\mathcal{M}}$
output : $\hat{\rho}$
$t=0$;
$\hat{\zeta}_{-n_{d}\Delta t}\sim\mathcal{GP}_{[\mathbf{y}_{\mathcal{M}}(-n_{d}\Delta t),\dots]},\quad x\in\mathcal{X}$;
repeat
       $\hat{\rho}(\cdot,t+\Delta t)\leftarrow\mathcal{G}_{n_{d}+1}[\hat{\zeta}_{t-n_{d}\Delta t}],\quad x\in\mathcal{X}$;
       $\hat{\zeta}_{t-n_{d}\Delta t+\Delta t}\leftarrow[\hat{\rho}(\cdot,t-n_{d}\Delta t+\Delta t),\dots]$;
       $t\leftarrow t+\Delta t$;
until end;
return $\hat{\rho}$
Algorithm 1 Open-Loop Observer $\hat{\mathcal{S}}^{ol}$
@@
While the purely autoregressive approach benefits from consistent predictions, it is sensitive to disturbance in the initial state. As reported in [21], autoregressive rollouts of neural operators suffer from unbounded out-of-distribution (OOD) error growth for long-term forecasts. Since we have access to online measurements, we compare $\hat{\mathcal{S}}^{ol}$ with another open-loop observer that resets the input state for every new prediction based on recent measurements. Specifically, the data-based state estimate $\hat{\zeta}^{\mathcal{D}}_{t}$ is obtained by interpolating the corresponding measurements $[\mathbf{y}_{\mathcal{M}}(t),\mathbf{y}_{\mathcal{M}}(t-\Delta t),\dots]$ using GP regression,
$\hat{\zeta}^{\mathcal{D}}_{t}\sim\mathcal{GP}_{[\mathbf{y}_{\mathcal{M}}(t),\mathbf{y}_{\mathcal{M}}(t-\Delta t),\dots]}\in\mathcal{H}^{n}_{\Omega}.$ (13)
@@
For comparability with (12), we consider a delayed state estimate $\hat{\zeta}^{\mathcal{D}}_{t-n_{d}\Delta t}$. The resulting dynamical system is denoted by an open-loop observer with reset $\hat{\mathcal{S}}^{ol\text{-}r}$, and defined as
@@
$\hat{\mathcal{S}}^{ol\text{-}r}[\mathcal{G},\mathbf{y}_{\mathcal{M}}]:\left\{\begin{array}[]{l}\hat{\rho}(\cdot,t+\Delta t)=\mathcal{G}_{n_{d}+1}[\hat{\zeta}^{\mathcal{D}}_{t-n_{d}\Delta t}],\\ \hfill t=0,\Delta t,\dots\\ \hat{\zeta}^{\mathcal{D}}_{t-n_{d}\Delta t}\sim\mathcal{GP}_{[\mathbf{y}_{\mathcal{M}}(t-n_{d}\Delta t),\dots]},\\ \hat{\zeta}^{\mathcal{D}}_{-n_{d}\Delta t}\sim\mathcal{GP}_{[\mathbf{y}_{\mathcal{M}}(-n_{d}\Delta t),\dots]}.\end{array}\right.$ (18)
The implementation details are given in Algorithm 2.
input : $\mathcal{G},\mathbf{y}_{\mathcal{M}}$
output : $\hat{\rho}$
$t=0$;
$\hat{\zeta}^{\mathcal{D}}_{-n_{d}\Delta t}\sim\mathcal{GP}_{[\mathbf{y}_{\mathcal{M}}(-n_{d}\Delta t),\dots]},\quad x\in\mathcal{X}$;
repeat
       $\hat{\rho}(\cdot,t+\Delta t)\leftarrow\mathcal{G}_{n_{d}+1}[\hat{\zeta}^{\mathcal{D}}_{t-n_{d}\Delta t}],\quad x\in\mathcal{X}$;
       $\hat{\zeta}^{\mathcal{D}}_{t-n_{d}\Delta t+\Delta t}\sim\mathcal{GP}_{[\mathbf{y}_{\mathcal{M}}(t-n_{d}\Delta t+\Delta t),\dots]}$;
       $t\leftarrow t+\Delta t$;
until end;
return $\hat{\rho}$
@@
Algorithm 2 Open-Loop Observer with Reset $\hat{\mathcal{S}}^{ol\text{-}r}$
@@
A limitation of $\hat{\mathcal{S}}^{ol\text{-}r}$ is that each prediction relies primarily on few recent data points. By property of the disconnected predictions, the observer does not leverage the full historical dataset.
@@ III-C Closed-Loop Observer
Combining the consistent predictions of $\hat{\mathcal{S}}^{ol}$ with the online updates of $\hat{\mathcal{S}}^{ol\text{-}r}$ leads to our main contribution: a closed-loop observer $\hat{\mathcal{S}}^{cl}$ that integrates an autoregressive rollout of $\mathcal{G}$ with online measurements using a correction operator $\mathcal{N}$. Analogous to a Kalman filter, $\mathcal{N}$ adjusts the predicted state estimate using recent measurements before it is passed as input to the next prediction. Measurements are introduced via a Luenberger-type error operator $\mathcal{E}$, which computes the error estimate $\hat{e}$ as the difference between the predicted state estimate $\hat{\zeta}$ and the data-based estimate $\hat{\zeta}^{\mathcal{D}}$. The correction operator $\mathcal{N}$ then processes $\hat{\zeta}$ and $\hat{e}$ to generate an updated state estimate $\hat{\zeta}^{u}\in\mathcal{H}^{n}_{\Omega}$, which is passed as input to the next prediction.
More precisely, we consider updates to a moving window of delayed extended state estimates,
$\begin{array}[]{l}\displaystyle\hat{\zeta}^{+}_{t-n_{d}\Delta t}=[\hat{\rho}(\cdot,t-(n-1)\Delta t),\dots,\\ \hfill\hat{\rho}(\cdot,t-n_{d}\Delta t),\dots,\\ \hfill\hat{\rho}(\cdot,t-n_{d}\Delta t-(n-1)\Delta t)],\\ \end{array}$ (19)
assuming $n_{d}>n-1$ as in the previous section. Thus, $\hat{\zeta}^{+}$ contains $\hat{\zeta}$ and extends forward in time such that the window length matches the prediction horizon (6) where $n_{out}=n_{d}+~{}1$ similar to (12) and (18). The correction is applied to this window of estimates, after which the updated state estimate $\hat{\zeta}^{u}$ is extracted as $\hat{\zeta}^{u+}_{-n:-1}$.
Thus, we define the Luenberger error operator $\mathcal{E}$ as
$\mathcal{E}[\hat{\zeta}^{+},\hat{\zeta}^{\mathcal{D}+}]\triangleq\hat{\zeta}^{+}-\hat{\zeta}^{\mathcal{D}+}\in\mathcal{H}^{n_{d}+1}_{\Omega}$
where $\hat{\zeta}^{\mathcal{D}+}$ is the extended data-based estimate. The correction operator $\mathcal{N}$ then processes the predicted state estimate $\hat{\zeta}^{+}$ and the error $\hat{e}^{+}=\mathcal{E}[\hat{\zeta}^{+},\hat{\zeta}^{\mathcal{D}+}]$ to generate the updated state estimate $\hat{\zeta}^{u+}$,
$\mathcal{N}:\begin{array}[t]{ccl}\mathcal{H}^{n_{d}+1}_{\Omega}\times\mathcal{H}^{n_{d}+1}_{\Omega}&\to&\mathcal{H}^{n_{d}+1}_{\Omega},\\ \hat{\zeta}^{+},\hat{e}^{+}&\mapsto&\hat{\zeta}^{u+}.\end{array}$ (20)
The correction update window $\hat{\zeta}^{+}_{t-n_{d}\Delta t}$ is shifted in increments of $\Delta t$ in parallel with the prediction input window $\hat{\zeta}_{t-n_{d}\Delta t}$, cf. (12). Since $\hat{\zeta}^{u}_{t-n_{d}\Delta t}\subset\hat{\zeta}^{u+}_{t-n_{d}\Delta t}$, the state estimates are updated before being passed as input to the next prediction. This yields the closed-loop observer $\hat{\mathcal{S}}^{cl}$,
@@
$\hat{\mathcal{S}}^{cl}[\mathcal{G},\mathcal{N},\mathbf{y}_{\mathcal{M}}]:\left\{\begin{array}[]{l}\hat{\rho}(\cdot,t+\Delta t)=\mathcal{G}_{n_{d}+1}[\hat{\zeta}^{u}_{t-n_{d}\Delta t}],\\ \hfill t=0,\Delta t,\dots\\ \hat{\zeta}^{u}_{t-n_{d}\Delta t}=\hat{\zeta}^{u+}_{-n:-1},\\ \hat{\zeta}^{u+}=\mathcal{N}\big{[}\hat{\zeta}^{+},\hat{e}^{+}\big{]},\\ \hat{e}^{+}=\mathcal{E}[\hat{\zeta}^{+},\hat{\zeta}^{\mathcal{D}+}],\\ \hat{\zeta}^{+}=[\hat{\rho}(\cdot,t-(n-1)\Delta t),\dots],\\ \hat{\zeta}^{\mathcal{D}+}\sim\mathcal{GP}_{[\mathbf{y}_{\mathcal{M}}(t-(n-1)\Delta t),\dots]},\\ \hat{\zeta}^{u}_{-n_{d}\Delta t}\sim\mathcal{GP}_{[\mathbf{y}_{\mathcal{M}}(-n_{d}\Delta t),\dots]}.\end{array}\right.$
The implementation details are outlined in Algorithm 3, and the system is illustrated in Fig. 2.
@@
$\hat{\zeta}^{u}_{-n_{d}\Delta t}\sim\mathcal{GP}_{[\mathbf{y}_{\mathcal{M}}(-n_{d}\Delta t),\dots]},\quad x\in\mathcal{X}$;
@@
       $\hat{\rho}(\cdot,t+\Delta t)\leftarrow\mathcal{G}_{n_{d}+1}[\hat{\zeta}^{u}_{t-n_{d}\Delta t}],\quad x\in\mathcal{X}$;
       $\hat{\zeta}^{+}\leftarrow[\hat{\rho}(\cdot,t-(n-1)\Delta t+\Delta t),\dots]$;
@@
       $\hat{\zeta}^{\mathcal{D}+}\sim\mathcal{GP}_{[\mathbf{y}_{\mathcal{M}}(t-(n-1)\Delta t+\Delta t),\dots]}$;
@@
       $\hat{e}^{+}=\mathcal{E}[\hat{\zeta}^{+},\hat{\zeta}^{\mathcal{D}+}]$;
       $\hat{\zeta}^{u+}\leftarrow\mathcal{N}\big{[}\hat{\zeta}^{+},\hat{e}^{+}\big{]}$;
@@
       $\hat{\zeta}^{u}_{t-n_{d}\Delta t+\Delta t}\leftarrow\hat{\zeta}^{u+}_{-n:-1}$;

### v2

@@ II-C Problem Statement
Following this discussion, we model traffic density flow as an $n^{\text{th}}$-order discrete-time dynamical system
$\mathcal{S}[\bar{\mathrm{P}}_{0}]:\begin{cases}\rho(\cdot,t+\Delta t)=\mathcal{G}[\mathrm{P}_{t}],\quad t>0,\\ \mathrm{P}_{t}=[\rho(\cdot,t),\rho(\cdot,t-\Delta t),\dots],\\ \mathrm{P}_{0}=\bar{\mathrm{P}}_{0},\end{cases}$ (6)
where $\rho(\cdot,t)\in\mathcal{H}^{1}_{\Omega}$ is the true density, $\mathrm{P}_{t}\in\mathcal{H}^{n}_{\Omega}$ is the state at time $t$, $\bar{\mathrm{P}}_{0}$ is the initial state, and $\mathcal{G}$ represents the unknown solution operator of the dynamics.
The objective of this paper is to develop an algorithm in the form of a dynamical system that observes (6) in closed-loop by generating real-time predictions $\hat{\rho}$ of the system state $\rho$ every $\Delta t$ time unit such that $\|\rho-\hat{\rho}\|$ is minimized. To achieve this goal, we assume access to online measurements $y(x,t)=\rho(x,t)+\epsilon$ with i.i.d. noise $\epsilon$ at stationary, sparse sensor locations $x\in\mathcal{M}\subset\Omega$, collectively denoted $\mathbf{y}_{\mathcal{M}}(t)$. We also assume access to offline observations $\{\mathcal{S}[\mathrm{P}_{0}^{i}]\}_{i\in\mathcal{I}}$, where $\mathrm{P}_{0}^{i}\sim\mu$ for some distribution $\mu$. The problem setting is illustrated in Fig. 1.
Figure 1: Problem setting. Our goal is to design an observer using offline data to learn the dynamics, and online measurements from stationary sensors at $x\in\mathcal{M}$ to calibrate predictions. The observer may then be used for macroscopic-level traffic control.
@@ III Methodology
To build a closed-loop observer of (6), the first step is to obtain a representation of the open-loop prediction, $\mathcal{G}$. Thus, our approach starts by learning an approximate solution operator $\mathcal{G}_{\theta}\approx\mathcal{G}$ from offline data. We then introduce two methods of iteratively applying $\mathcal{G}$ to observe (6) in an open-loop manner, $\hat{\mathcal{S}}^{ol}$ and $\hat{\mathcal{S}}^{ol\text{-}r}$. To close the loop, we finally design an algorithm, $\hat{\mathcal{S}}^{cl}$, that integrates sparse online measurements with recursive predictions using $\mathcal{G}$. The research problem is then addressed by using $\hat{\mathcal{S}}^{cl}$ together with $\mathcal{G}_{\theta}$, resulting in a data-driven, closed-loop observer.
@@ III-A Solution Operator: Model Identification
Given the offline scenarios $\{\mathcal{S}[\mathrm{P}_{0}^{i}]\}_{i\in\mathcal{I}}$, the goal is to learn a representation of the solution operator
$\mathcal{G}_{k}:\begin{array}[t]{rcl}\mathcal{H}^{n}_{\Omega}&\to&\mathcal{H}^{1}_{\Omega},\\ \mathrm{P}_{0}&\mapsto&\rho_{k\Delta t},\end{array}$
which maps the input state $\mathrm{P}_{0}$ to the solution function at $k$ time steps later, $\rho_{k\Delta t}$. While both $\mathrm{P}_{0}$ and $\rho_{k\Delta t}$ are infinite-dimensional functions, we assume that the available data is represented in finite-dimensional form, obtained by evaluating them on a uniformly discretized grid $\mathcal{X}\subset\Omega$. This yields $\mathrm{P}_{0}\rvert_{\mathcal{X}}\in\mathbb{R}^{n\times\lvert\mathcal{X}\rvert}$ and $\rho_{k\Delta t}\rvert_{\mathcal{X}}\in\mathbb{R}^{\lvert\mathcal{X}\rvert}$. To bridge the gap between finite-dimensional data and function space learning, the approximate operator $\mathcal{G}_{\theta}$ is modeled by an FNO (2)–(4).
@@
$\mathcal{G}_{\theta}:\begin{array}[t]{rcl}\mathcal{H}^{n}_{\Omega}&\to&\mathcal{H}^{n_{out}}_{\Omega},\\ \mathrm{P}_{0}&\mapsto&[\hat{\rho}_{\Delta t},\hat{\rho}_{2\Delta t},\dots,\hat{\rho}_{n_{out}\Delta t}].\end{array}$ (7)
Including the horizon up to the target estimate $\hat{\rho}_{n_{out}\Delta t}$ was empirically found to improve performance, and it reflects the setup in [25]. Thus, the corresponding data-driven learning objective [^2] While we focus on a data-driven objective in this paper, we highlight that it is possible to regularize using lower-order models such as (1) with a predefined or jointly learned flow function $Q$ following [24, 16]. is given by
$\min_{\theta}\mathbb{E}_{a\sim\mu}\bigg{[}\sum\limits_{k=1}^{n_{out}}\int\limits_{\Omega}\|\mathcal{G}_{k}[a](x)-\mathcal{G}_{\theta}^{(k)}[a](x)\|^{2}dx\bigg{]}.$ (8)
Fig. 2 illustrates an example forward pass of a learned operator $\mathcal{G}_{\theta}$ for $n_{out}=100$. We refer to this as open-loop prediction in the observer framework.
Figure 2: Open-loop prediction with solution operator $\mathcal{G}_{\theta}$. It predicts $\hat{\rho}_{\Delta t},\hat{\rho}_{2\Delta t},...,\hat{\rho}_{n_{out}\Delta t}$ given an initial state $\mathrm{P}_{0}=[\rho_{0},\rho_{-\Delta t},\dots,\rho_{-(n-1)\Delta t}]$. Here, $n=10$ and $n_{out}=100$.
@@ III-B Open-Loop Observers
$\hat{\mathrm{P}}_{t}=[\hat{\rho}(\cdot,t),\hat{\rho}(\cdot,t-\Delta t),\dots]\in\mathcal{H}^{n}_{\Omega}.$
Thus, triggering an autoregressive rollout only requires an estimate of the initial state $\hat{\mathrm{P}}_{0}$. However, this is only partially observed in our problem, since measurements are restricted to the sparse sensor locations $\mathcal{M}\subsetneq\Omega$. To complete the initial state estimate, we interpolate between the sparse measurements $\mathbf{y}_{\mathcal{M}}(t)$ with GP regression,
$\hat{\mathrm{P}}_{0}\sim\mathcal{GP}_{[\mathbf{y}_{\mathcal{M}}(0),\mathbf{y}_{\mathcal{M}}(-\Delta t),\dots]}\in\mathcal{H}^{n}_{\Omega}.$
@@
We found that using multi-step ahead prediction $\mathcal{G}_{n_{d}+1}$ with a delayed state estimate $\hat{\mathrm{P}}_{t-n_{d}\Delta t}$ offset by $n_{d}>n-1$ improved stability. This creates the open-loop observer $\hat{\mathcal{S}}^{ol}$:
@@
$\hat{\mathcal{S}}^{ol}[\mathcal{G},\mathbf{y}_{\mathcal{M}}]:\left\{\begin{array}[]{l}\hat{\rho}(\cdot,t+\Delta t)=\mathcal{G}_{n_{d}+1}[\hat{\mathrm{P}}_{t-n_{d}\Delta t}],\\ \hfill t=0,\Delta t,\dots\\ \hat{\mathrm{P}}_{t-n_{d}\Delta t}=[\hat{\rho}(\cdot,t-n_{d}\Delta t),\dots],\\ \hat{\mathrm{P}}_{-n_{d}\Delta t}\sim\mathcal{GP}_{[\mathbf{y}_{\mathcal{M}}(-n_{d}\Delta t),\dots]}.\end{array}\right$. (13)
@@
The gap between the initial state and the first estimate, $\hat{\rho}(\cdot,t),$ $t\in[-n_{d}\Delta t+\Delta t,\dots,0]$, is filled by interpolating the corresponding measurements with GP regression. We illustrate the observer in Fig. 3.
Figure 3: Open-loop observer $\hat{\mathcal{S}}^{ol}$
@@
While the purely autoregressive approach benefits from consistent predictions, it is sensitive to disturbance in the initial state. As reported in [26], autoregressive rollouts of neural operators suffer from unbounded out-of-distribution (OOD) error growth for long-term forecasts. Since we have access to online measurements, we compare $\hat{\mathcal{S}}^{ol}$ with another open-loop observer that resets the input state for every new prediction based on recent measurements. Specifically, the data-based state estimate $\hat{\mathrm{P}}^{\mathcal{D}}_{t}$ is obtained by interpolating the corresponding measurements $\mathbf{y}_{\mathcal{M}}(t),\mathbf{y}_{\mathcal{M}}(t-\Delta t),\dots]$ using GP regression,
$\hat{\mathrm{P}}^{\mathcal{D}}_{t}\sim\mathcal{GP}_{[\mathbf{y}_{\mathcal{M}}(t),\mathbf{y}_{\mathcal{M}}(t-\Delta t),\dots]}\in\mathcal{H}^{n}_{\Omega}.$ (14)
@@
For comparability with (13), we consider a delayed state estimate $\hat{\mathrm{P}}^{\mathcal{D}}_{t-n_{d}\Delta t}$. The resulting dynamical system is denoted open-loop observer with reset $\hat{\mathcal{S}}^{ol\text{-}r}$, and defined as
@@
$\hat{\mathcal{S}}^{ol\text{-}r}[\mathcal{G},\mathbf{y}_{\mathcal{M}}]:\left\{\begin{array}[]{l}\hat{\rho}(\cdot,t+\Delta t)=\mathcal{G}_{n_{d}+1}[\hat{\mathrm{P}}^{\mathcal{D}}_{t-n_{d}\Delta t}],\\ \hfill t=0,\Delta t,\dots\\ \hat{\mathrm{P}}^{\mathcal{D}}_{t-n_{d}\Delta t}\sim\mathcal{GP}_{[\mathbf{y}_{\mathcal{M}}(t-n_{d}\Delta t),\dots]},\\ \hat{\mathrm{P}}^{\mathcal{D}}_{-n_{d}\Delta t}\sim\mathcal{GP}_{[\mathbf{y}_{\mathcal{M}}(-n_{d}\Delta t),\dots]}.\end{array}\right$. (19)
The observer is illustrated in Fig. 4.
@@
Figure 4: Open-loop observer with reset $\hat{\mathcal{S}}^{ol\text{-}r}$
@@
A limitation of $\hat{\mathcal{S}}^{ol\text{-}r}$ is that each prediction relies on a few data points. By the property of the disconnected predictions, the observer does not leverage the full historical dataset.
@@ III-C Closed-Loop Observer
Combining the consistent predictions of $\hat{\mathcal{S}}^{ol}$ with the online updates of $\hat{\mathcal{S}}^{ol\text{-}r}$ leads to our main contribution: a closed-loop observer $\hat{\mathcal{S}}^{cl}$ that integrates an autoregressive rollout of $\mathcal{G}$ with online measurements using a correction operator $\mathcal{N}$. Analogous to a Kalman filter, $\mathcal{N}$ adjusts the predicted state estimate using recent measurements before it is passed as input to the next prediction. Measurements are introduced via a Luenberger-type error operator $\mathcal{E}$, which computes the error estimate $\hat{e}$ as the difference between the predicted state estimate $\hat{\mathrm{P}}$ and the data-based estimate $\hat{\mathrm{P}}^{\mathcal{D}}$. The correction operator $\mathcal{N}$ then processes $\hat{\mathrm{P}}$ and $\hat{e}$ to generate an updated state estimate $\hat{\mathrm{P}}^{u}\in\mathcal{H}^{n}_{\Omega}$, which is passed as input to the next prediction. We summarize $\hat{\mathcal{S}}^{cl}$ as
@@
$\hskip-5.69046pt\hat{\mathcal{S}}^{cl}[\mathcal{G},\mathcal{N},\mathbf{y}_{\mathcal{M}}]:\left\{\begin{array}[]{l}\hat{\rho}(\cdot,t+\Delta t)=\mathcal{G}_{n_{d}+1}[\hat{\mathrm{P}}^{u}_{t-n_{d}\Delta t}],\\ \hfill t=0,\Delta t,\dots\\ \hat{\mathrm{P}}^{u}_{t-n_{d}\Delta t}=\mathcal{N}\big{[}\hat{\mathrm{P}},\hat{e}\big{]},\\ \hat{e}=\mathcal{E}[\hat{\mathrm{P}},\hat{\mathrm{P}}^{\mathcal{D}}],\\ \hat{\mathrm{P}}_{t-n_{d}\Delta t}=[\hat{\rho}(\cdot,t-n_{d}\Delta t),\dots],\\ \hat{\mathrm{P}}^{\mathcal{D}}_{t-n_{d}\Delta t}\sim\mathcal{GP}_{[\mathbf{y}_{\mathcal{M}}(t-n_{d}\Delta t),\dots]},\\ \hat{\mathrm{P}}^{u}_{-n_{d}\Delta t}\sim\mathcal{GP}_{[\mathbf{y}_{\mathcal{M}}(-n_{d}\Delta t),\dots]},\end{array}\right$. (27)
and illustrate the system in Fig. 5.
Figure 5: Closed-loop observer $\hat{\mathcal{S}}^{cl}$
In implementing (27), corrections are applied to a moving window of historical predictions,
[…]
assuming $n_{d}>n-1$ as in the previous section. Thus, the extended state estimate $\hat{\mathrm{P}}^{+}$ contains $\hat{\mathrm{P}}$ and extends forward in time such that the window length matches the prediction horizon (7) where $n_{out}=n_{d}+~1$ similar to (13) and (19). After applying the correction to this window, the updated state estimate $\hat{\mathrm{P}}^{u}$ is taken as the subset $\hat{\mathrm{P}}^{u+}_{-n:-1}$. The procedure is illustrated in Fig. 6.
Figure 6: Implementation of $\hat{\mathcal{S}}^{cl}$ with extended state estimates
With $\hat{\mathrm{P}}^{+}$ and the corresponding extended data-based estimate $\hat{\mathrm{P}}^{\mathcal{D}+}$, the Luenberger error operator $\mathcal{E}$ is defined as
$\mathcal{E}[\hat{\mathrm{P}}^{+},\hat{\mathrm{P}}^{\mathcal{D}+}]\triangleq\hat{\mathrm{P}}^{+}-\hat{\mathrm{P}}^{\mathcal{D}+}\in\mathcal{H}^{n_{d}+1}_{\Omega}.$
The correction operator $\mathcal{N}$ processes $\hat{\mathrm{P}}^{+}$ and the error $\hat{e}^{+}=\mathcal{E}[\hat{\mathrm{P}}^{+},\hat{\mathrm{P}}^{\mathcal{D}+}]$ to generate the updated state estimate $\hat{\mathrm{P}}^{u+}$,
$\mathcal{N}:\begin{array}[t]{ccl}\mathcal{H}^{n_{d}+1}_{\Omega}\times\mathcal{H}^{n_{d}+1}_{\Omega}&\to&\mathcal{H}^{n_{d}+1}_{\Omega},\\ \hat{\mathrm{P}}^{+},\hat{e}^{+}&\mapsto&\hat{\mathrm{P}}^{u+}.\end{array}$ (29)
The correction update window $\hat{\mathrm{P}}^{+}_{t-n_{d}\Delta t}$ is shifted in increments of $\Delta t$ in parallel with the prediction input window $\hat{\mathrm{P}}_{t-n_{d}\Delta t}$, cf. (13). Since $\hat{\mathrm{P}}^{u}_{t-n_{d}\Delta t}\subset\hat{\mathrm{P}}^{u+}_{t-n_{d}\Delta t}$, the state estimates are updated before being passed as input to the next prediction. The implementation details are outlined in Algorithm 1.
@@
$\hat{\mathrm{P}}^{u}_{-n_{d}\Delta t}\sim\mathcal{GP}_{[\mathbf{y}_{\mathcal{M}}(-n_{d}\Delta t),\dots]},\quad x\in\mathcal{X}$;
@@
 $\hat{\rho}(\cdot,t+\Delta t)\leftarrow\mathcal{G}_{n_{d}+1}[\hat{\mathrm{P}}^{u}_{t-n_{d}\Delta t}],\quad x\in\mathcal{X}$;
 $\hat{\mathrm{P}}^{+}\leftarrow[\hat{\rho}(\cdot,t-(n-1)\Delta t+\Delta t),\dots]$;
@@
 $\hat{\mathrm{P}}^{\mathcal{D}+}\sim\mathcal{GP}_{[\mathbf{y}_{\mathcal{M}}(t-(n-1)\Delta t+\Delta t),\dots]}$;
@@
 $\hat{e}^{+}=\mathcal{E}[\hat{\mathrm{P}}^{+},\hat{\mathrm{P}}^{\mathcal{D}+}]$;
 $\hat{\mathrm{P}}^{u+}\leftarrow\mathcal{N}\big{[}\hat{\mathrm{P}}^{+},\hat{e}^{+}\big{]}$;
@@
 $\hat{\mathrm{P}}^{u}_{t-n_{d}\Delta t+\Delta t}\leftarrow\hat{\mathrm{P}}^{u+}_{-n:-1}$;